OpenStudy (anonymous):

(02.01) Solve for x: 3/4x + 5/8= 4x

OpenStudy (anonymous):

3/4x + 5/8= 4x Get the x's on the same side 5/8 = 4x -3/4x simplify and solve for x

OpenStudy (anonymous):

5/8 = 16/3x To get x by itself multiply by the reciprocal of 16/3, which is 3/16 5/8 * 3/16 = x 15/128=x

OpenStudy (anonymous):

Thank you so much for your help

OpenStudy (anonymous):

you're welcome